# Beatty solutions of almost Golomb equations

### 16 Apr 2026 v2

@@ 4.5 Wythoff endpoints
Let $\alpha=\sqrt{2}$.
  1. 1.
    At $d=\alpha/2$, $a(n)=\lfloor n/\alpha+\alpha/2\rfloor$ and the complement $\bar{a}(n)=\lfloor n\alpha+\alpha/2\rfloor$ form a Beatty partition of $\mathbb{Z}_{\geq 1}$.
  2. 2.
    At $d=2(\alpha{-}1)$, $a(n)=\lfloor cn+2(\alpha{-}1)\rfloor$ and $\bar{a}(n)=n+a(n)$ form a complementary pair in the $\alpha$-Wythoff game.
The lower endpoint gives the standard Beatty partition via Rayleigh’s theorem (since $1/\alpha+1/(1+\alpha)=1$). The upper endpoint is the $\alpha$-Wythoff pair studied by Fraenkel [4].
@@
At $d=\alpha/2$, the canonical solution satisfies both the strong identity $a(f(n))=n$ and the Beatty complementation. This double constraint pins down the shift uniquely. At the upper endpoint $d=2(\alpha{-}1)$, only the triple-nested identity holds, and the Beatty complementation is replaced by the Wythoff partition.
@@ 6 The defect at the right endpoint
6.1 Definition
For $d=2(\sqrt{2}{-}1)$ (the right endpoint of $I$), the triple-nested identity $a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n)$ holds by Theorem 3. The strong identity $a_{\mathrm{B}}(f(n))=n$ however fails at certain positions. Define the defect
@@ 6.1 Definition
and the defect set $\mathcal{D}=\{n\geq 1:\delta(n)=1\}$. The first elements of $\mathcal{D}$ are
@@
6.2 The absorption theorem
The central result of this section is that the defect positions are characterised by a simple interval condition on the Sturmian address $\theta_{n-1}$, and in particular every defect position is a repeat of $a_{\mathrm{B}}$.
@@ 6.2 The absorption theorem
Applying the floor, $a_{\mathrm{B}}(f(n))=n+1$ precisely when $n{+}1\leq cf(n)+d<n+2$, which using (19) rearranges to $\tfrac{1}{2}\leq-c(\theta_{n}+\theta_{n-1})<\tfrac{1}{2}$, equivalently $\theta_{n}+\theta_{n-1}\leq 1/(2c)=c$ (the lower bound is automatic).
@@
Case $\theta_{n-1}<1-c$. Then $\theta_{n}=\theta_{n-1}+c$, and $\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c\leq c$ forces $\theta_{n-1}=0$, which never happens since $d$ is irrational.
Case $\theta_{n-1}\geq 1-c$. Then $\theta_{n}=\theta_{n-1}+c-1$, and $\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c-1\leq c$ becomes $\theta_{n-1}\leq 1/2$, yielding $\theta_{n-1}\in[1-c,1/2]$.
@@
The absorption mechanism is now transparent: at a defect position, the Sturmian address $\theta_{n}$ lies in the low interval [0,c1/2][0,c-1/2], forcing $a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$. The identity (5) then survives because the extra +1+1 in $a_{\mathrm{B}}(f(n))$ is absorbed: $a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$.
@@ 6.3 The gap sequence
The characterisation $\mathcal{D}=\{n\geq 2:\theta_{n-1}\in[1-c,1/2]\}$ expresses $\mathcal{D}$ as the return-time set of the rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1-c,1/2]. By the three-gap theorem (Świerczkowski [10]), the gaps take at most three values. Numerical experiments confirm these three values are $3$, 44, and $7$, with empirical frequencies
$\sqrt{2}-1,\quad 3-2\sqrt{2},\quad\sqrt{2}-1,$
and average gap $3(\sqrt{2}{-}1)+4(3{-}2\sqrt{2})+7(\sqrt{2}{-}1)=2(\sqrt{2}{+}1)$. The corresponding density $1/(2(\sqrt{2}{+}1))=(\sqrt{2}{-}1)/2$ matches Theorem 20(b).
The letter $4$ is always preceded and followed by $3$: every $4$ appears in a block $3,4,3$ (verified for $k\leq 10^{5}$).
@@
The gap sequence of $\mathcal{D}$ on the alphabet $\{3,4,7\}$ is a coding of the fixed point of a primitive substitution; equivalently, $\mathcal{D}$ is a morphic set.

### 21 Apr 2026 v3

@@ 4.6 The triple-nested identity for $r=3$
The sequence $a(n)=\lfloor cn+d\rfloor$ with $c=1/\!\sqrt{3}$ satisfies $a(a(f(n)))=a(n)$ for all $n\geq 3$ if and only if
$d\in I_{3}=\Bigl[\,\frac{\sqrt{3}+3}{6},\;\frac{5\sqrt{3}-3}{6}\,\Bigr].$ (27)
The interval is centred at $\sqrt{3}/2$ with half-width $(2\sqrt{3}-3)/6$, and both endpoints lie in $\mathbb{Q}(\sqrt{3})$.
The proof occupies this subsection. We describe the four combinatorial regimes, prove sufficiency, then prove necessity by accumulating the obstruction at $\theta=1-c$.
The four regimes
Parametrise by $t=\theta_{n-2}\in[0,1)$, so that $\theta_{n-1}=\{t+c\}$ and $\theta_{n}=\{t+2c\}$. Since $c\in(1/2,1)$ and $2c\in(1,2)$, the wraps occur at t=1ct=1-c for $\theta_{n-1}$ and at t=22ct=2-2c and t=12c(<0)t=1-2c\,(<0) for $\theta_{n}$. A direct case analysis against the threshold $\theta_{\cdot}<1-c$ that defines a repeat yields the four regimes
[…]
The sum $\Theta(n)=t+\theta_{n-1}+\theta_{n}$ is piecewise affine in $t$ and takes the values
$\Theta(t)=\begin{cases}3t+3c-1,&t\in\mathrm{A}\cup\mathrm{B},\\[2.0pt] 3t+3c-2,&t\in\mathrm{C},\\[2.0pt] 3t+3c-3,&t\in\mathrm{D}.\end{cases}$ (28)
The ranges of $\Theta$ on the four regimes are
$\Theta(\mathrm{A})$ $=[\sqrt{3}-1,5-2\sqrt{3}),$ (29)
$\Theta(\mathrm{B})$ $=[5-2\sqrt{3},2),$
$\Theta(\mathrm{C})$ $=[1,4-\sqrt{3}),$
$\Theta(\mathrm{D})$ $=[3-\sqrt{3},\;\sqrt{3}).$
Two facts drive the proof.
Fact A (lower extremum). The non-repeat positions for $n$ are exactly $\mathrm{B}\cup\mathrm{C}$, and
$\inf_{t\in\mathrm{B}\cup\mathrm{C}}\Theta(t)=1,$
attained only at the left endpoint t=1ct=1-c of regime C.
Fact B (upper extremum). The non-repeat positions for n1n-1 are exactly $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$, and
$\sup_{t\in\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}}\Theta(t)=2,$
approached only at the left limit $t\to(1-c)^{-}$ from regime B.
Both facts follow from (29) by direct inspection.
Two key threshold values for $D$
Compute $D$ at the two endpoints of $I_{3}$. At $d_{L}=(\sqrt{3}+3)/6$:
$D_{L}=\frac{\sqrt{3}+3}{6}\cdot(\sqrt{3}+1)=\frac{(\sqrt{3}+3)(\sqrt{3}+1)}{6}=\frac{3+\sqrt{3}+3\sqrt{3}+3}{6}=\frac{6+4\sqrt{3}}{6}=1+\frac{2}{\sqrt{3}}=1+2c.$
At $d_{R}=(5\sqrt{3}-3)/6$:
$D_{R}=\frac{5\sqrt{3}-3}{6}\cdot(\sqrt{3}+1)=\frac{15+5\sqrt{3}-3\sqrt{3}-3}{6}=\frac{12+2\sqrt{3}}{6}=2+\frac{1}{\sqrt{3}}=2+c.$
So $d\in I_{3}$ is equivalent to $D\in[1+2c,\,2+c]$.
Sufficiency
Assume $d\in I_{3}$, equivalently $D\in[1+2c,\,2+c]$. We first check that the shift $m(n)$ is bounded a priori, so that (N+) and (N-) together become sufficient. The sum $\Theta(n)$ ranges over $\Theta(\mathrm{A})\cup\Theta(\mathrm{B})\cup\Theta(\mathrm{C})\cup\Theta(\mathrm{D})\subset[\sqrt{3}-1,\,4-\sqrt{3})$ by (29). Multiplying by $c$ and using $c\sqrt{3}=1$,
$c\Theta(n)\in[1-c,\,4c-1),$
whence
$D-c\Theta(n)\in(D-4c+1,\,D-1+c]\subset(1+2c-4c+1,\,2+c-1+c]=(2-2c,\,1+2c].$
In particular $D-c\Theta(n)\in(0,3)$, so $\lfloor D-c\Theta(n)\rfloor\in\{0,1,2\}$ and by (26)
$m(n)\in\{-1,0,1\}\qquad\text{for every }n\geq 3,\ d\in I_{3}.$ (30)
With (30) at hand, (N+) and (N-) jointly imply the triple-nested identity on $I_{3}$. We now verify them regime by regime.
Requirement (N+) forbids $m(n)\geq 1$ at non-repeats of $n$, which by (26) is the inequality $D-c\Theta(n)\geq 2$, equivalently
$\Theta(n)\leq\frac{D-2}{c}=\sqrt{3}(D-2).$ (31)
The hypothesis $D\leq 2+c$ gives $\sqrt{3}(D-2)\leq\sqrt{3}c=1$.
For $D<2+c$ (the interior of $I_{3}$ on the right), the bound in (31) is strict, $\sqrt{3}(D-2)<1$, and Fact A gives $\Theta(n)\geq 1>\sqrt{3}(D-2)$ on $\mathrm{B}\cup\mathrm{C}$. Hence (N+) holds vacuously.
For $D=2+c$ (the right endpoint $d=d_{R}$), equality in Fact A is possible only at the single point t=1ct=1-c. Writing $d_{R}=(5c-1)/2$, the equation $\{c(n-2)+d_{R}\}=1-c$ unfolds to $c(n-1)+(5c-1)/2-1\in\mathbb{Z}$. Multiplying by $2$, this reads $(2n+3)c=2k+3$ for some integer $k$. Since $2n+3\neq 0$ and $c$ is irrational, no integer $n$ solves this equation, the threshold is never attained, and (N+) holds.
Requirement (N-) forbids $m(n)\leq-1$ at non-repeats of n1n-1, which by (26) is the inequality $D-c\Theta(n)<1$, equivalently
$\Theta(n)>\frac{D-1}{c}=\sqrt{3}(D-1).$ (32)
The hypothesis $D\geq 1+2c$ gives $\sqrt{3}(D-1)\geq 2\sqrt{3}c=2$, and by Fact B $\Theta(n)\leq 2$ on $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$ with the supremum not attained. For $D>1+2c$ the bound is strict and (N-) holds vacuously. At $D=1+2c$ (the left endpoint $d=d_{L}=(1+c)/2$), the bound becomes $\Theta(n)>2$, which is impossible on $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$ since the supremum $2$ is not attained. Hence (N-) holds.
Sufficiency is established for every $d\in I_{3}$.
Necessity
Right bound. Assume $d>d_{R}$, equivalently $D>2+c$. Then $\sqrt{3}(D-2)>1$, so the threshold $\sqrt{3}(D-2)$ lies strictly inside the range $\Theta(\mathrm{C})=[1,4-\sqrt{3})$, exceeding the left endpoint. The set
$U_{R}=\{t\in\mathrm{C}:\Theta(t)<\sqrt{3}(D-2)\}=\{t:1-c\leq t<(1-c)+(D-2-c)/\sqrt{3}\}$
is a non-empty open sub-interval of regime C accumulating at t=1ct=1-c. On $U_{R}$, we have $D-c\Theta(n)>2$, hence $m(n)\geq 1$, while $n$ is a non-repeat (regime C). This violates (N+). The orbit $(\theta_{n-2})_{n\geq 3}$ is dense in $[0,1)$ by Weyl’s theorem, so it visits $U_{R}$ infinitely often, and (5) fails at infinitely many $n$.
Left bound. Assume $d<d_{L}$, equivalently $D<1+2c$. Then $\sqrt{3}(D-1)<2$, so the threshold $\sqrt{3}(D-1)$ lies strictly below the supremum $2$ on $\Theta(\mathrm{B})=[5-2\sqrt{3},2)$. The set
$U_{L}=\{t\in\mathrm{B}:\Theta(t)>\sqrt{3}(D-1)\}=\{t:t_{\star}<t<1-c\},$
where $t_{\star}=\max\bigl((\sqrt{3}(D-1)+1)/3-c,\,2-3c\bigr)$, is a non-empty open sub-interval of regime B accumulating at the left limit of t=1ct=1-c. On $U_{L}$, we have $D-c\Theta(n)<1$, hence $m(n)\leq-1$, while n1n-1 is a non-repeat (regime B). This violates (N-). Density of the orbit yields infinitely many failures.
Endpoints. At $d=d_{R}$, $D=2+c$ and the threshold for $m(n)\geq 1$ is $\Theta\leq 1$. The only value in $\Theta(\mathrm{B}\cup\mathrm{C})$ satisfying this is $\Theta=1$ at t=1ct=1-c, never attained. At $d=d_{L}$, $D=1+2c$ and the threshold for $m(n)\leq-1$ is $\Theta>2$, never attained on $\mathrm{A}\cup\mathrm{B}\cup\mathrm{D}$. Both endpoints are therefore included in $I_{3}$.
This completes the proof of Theorem 21.
@@
The interval $I_{3}$ is symmetric around the strong-identity shift $\sqrt{3}/2$, with half-width $(2\sqrt{3}-3)/6$. Symmetry for $r=3$ contrasts with the asymmetry of $I$ for $r=2$. In both cases, the mechanism is the same. The threshold for $m(n)=+1$ and the threshold for m(n)=1m(n)=-1 are governed by extrema of $\Theta$ over the non-repeat sets of $n$ and of n1n-1 respectively, and for $r=3$ these two extrema sit symmetrically about the centre. For $r=2$, the two-term window gives only one non-repeat regime coupling both constraints, producing asymmetry.
Remark 23.
The same method applies to any non-square $r\geq 2$ with $c=1/\!\sqrt{r}$ and $f(n)=\sum_{j=0}^{r-1}a(n{-}j)$. The orbit $(\theta_{n-r+1},\dots,\theta_{n})$ partitions $[0,1)$ into finitely many regimes labelled by the joint repeat or non-repeat status of (n,n1)(n,n-1). On each regime the combined fractional-part sum $\Theta(n)$ is piecewise affine in the base parameter $t=\theta_{n-r+1}$. Local thresholds of $D-c\Theta(n)$ against the integers produce both the sufficiency bounds and the failure sets, and orbit density yields infinitely many failures outside the resulting interval $I_{r}$. We expect $I_{r}$ to be centred at $\sqrt{r}/2$, symmetric for odd $r$ and asymmetric for even $r$.
At the right endpoint $d=(5\sqrt{3}{-}3)/6$, the defect set begins
[…]
(A395252), with gap values $\{7,\,12,\,19\}$ where 19=7+1219=7{+}12. The three values correspond to consecutive convergent denominators of $1/\!\sqrt{3}=[0;1,1,2,1,2,1,2,\ldots]$, and the density of the defect set in $\mathbb{Z}_{\geq 1}$ equals $(2{-}\sqrt{3})/3$, matching the length of the rotation interval controlling the absorption. A substitution analogue of Conjecture 26 for $r=3$ is formulated as an open problem in Section 7.
@@ 6 The defect at the right endpoint
At the right endpoint $d=2(\sqrt{2}{-}1)$ of the interval $I$, the triple-nested identity holds but the strong identity $a_{\mathrm{B}}(f(n))=n$ fails at a sparse set of positions which we call the defect set $\mathcal{D}$. This section describes the structure of $\mathcal{D}$ in three steps. First we characterise $\mathcal{D}$ as the set of return times of the Sturmian rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1-c,1/2] (Theorem 24), which explains why every defect position is a repeat of $a_{\mathrm{B}}$, so that the extra increment in $a_{\mathrm{B}}(f(n))$ is absorbed by the outer $a$. Second, the three-distance theorem gives three possible gap values between consecutive defect positions, namely $3$, 44, and $7$. Third, the gap sequence on the alphabet $\{3,4,7\}$ is conjecturally generated by an explicit primitive substitution with Perron eigenvalue $1+\sqrt{2}$ (Conjecture 26), tying the defect structure to the Pell recurrence.
For $d=2(\sqrt{2}{-}1)$, the triple-nested identity $a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n)$ holds by Theorem 3, and expansion of the defining formula shows $a_{\mathrm{B}}(f(n))\in\{n,\,n{+}1\}$. Define the defect
@@
and the defect set $\mathcal{D}=\{n\geq 1:\delta(n)=1\}$. The first elements of $\mathcal{D}$ (A395251) are
@@
6.1 Defect positions as return times
We now show that the defect positions are exactly the return times of the Sturmian rotation to an explicit interval, and that every defect position is a repeat of $a_{\mathrm{B}}$.
@@ 6.1 Defect positions as return times
Applying the floor, $a_{\mathrm{B}}(f(n))=n+1$ precisely when $n{+}1\leq cf(n)+d<n+2$, which using (33) rearranges to
$\tfrac{1}{2}\leq-c(\theta_{n}+\theta_{n-1})<\tfrac{1}{2},$
equivalently $\theta_{n}+\theta_{n-1}\leq 1/(2c)=c$ (the lower bound is automatic).
@@
If $\theta_{n-1}<1-c$, then $\theta_{n}=\theta_{n-1}+c$, and
$\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c\leq c$
forces $\theta_{n-1}=0$, which never happens since $d$ is irrational.
If $\theta_{n-1}\geq 1-c$, then $\theta_{n}=\theta_{n-1}+c-1$, and
$\theta_{n}+\theta_{n-1}=2\theta_{n-1}+c-1\leq c$
becomes $\theta_{n-1}\leq 1/2$, yielding $\theta_{n-1}\in[1-c,1/2]$.
@@
The absorption mechanism now reads explicitly. At a defect position, the Sturmian address $\theta_{n}$ lies in the low interval [0,c1/2][0,c-1/2], forcing $a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n)$. The identity (5) then survives because the extra +1+1 in $a_{\mathrm{B}}(f(n))$ is absorbed by the outer application of $a_{\mathrm{B}}$,
$a_{\mathrm{B}}(a_{\mathrm{B}}(f(n)))=a_{\mathrm{B}}(n{+}1)=a_{\mathrm{B}}(n).$
@@ 6.2 The gap sequence
The characterisation $\mathcal{D}=\{n\geq 2:\theta_{n-1}\in[1-c,1/2]\}$ expresses $\mathcal{D}$ as the return-time set of the rotation $\theta\mapsto\{\theta+c\}$ to the interval [1c,1/2][1-c,1/2]. By the three-distance theorem of Sós [11] (see Alessandri and Berthé [10] for a modern exposition within combinatorics on words), the gaps take at most three values. The return-time analysis of Proposition 25 below identifies these values as $3$, 44, and $7$, and computes the corresponding letter frequencies.
The letter $4$ is always preceded and followed by $3$. Every occurrence of $4$ in the gap sequence appears within a block $3,4,3$, verified numerically for $k\leq 10^{5}$.
Proposition 25.
The gap sequence of $\mathcal{D}$ on the alphabet $\{3,4,7\}$ has letter frequencies
$f_{3}=\sqrt{2}-1,\quad f_{4}=3-2\sqrt{2},\quad f_{7}=\sqrt{2}-1,$
matching Theorem 24(b), and the average gap is $2(\sqrt{2}+1)$. The gap sequence of $\mathcal{D}$ is A395253.
Proof.
The defect set is $\mathcal{D}=\{n\geq 2:\theta_{n-1}\in J\}$ with J=[1c, 1/2]J=[1-c,\,1/2], so $|J|=c-1/2=(\sqrt{2}-1)/2$. For each return time $i\geq 1$, let
$J_{i}=\{\theta\in J:R_{c}^{i}\theta\in J\text{ and }R_{c}^{k}\theta\notin J\text{ for }1\leq k<i\},$
where $R_{c}\colon\theta\mapsto\{\theta+c\}$. We compute the sub-intervals directly from the iterates. Using $c\in(1/2,1)$ and 2c<3c1<3c<22c<3c-1<3c<2, the values $kc\bmod 1$ for small $k$ are
$R_{c}\theta=\theta+c-1,\qquad R_{c}^{2}\theta=\theta+2c-1,\qquad R_{c}^{3}\theta=\theta+3c-2$
for $\theta\in J$. Neither $R_{c}\theta\in[0,c-1/2]$ nor $R_{c}^{2}\theta\in[c,2c-1/2]$ meets $J$, so every return time is at least $3$.
Return time $3$ occurs when $\theta+3c-2\in[1-c,1/2]$, equivalently $\theta\in[3-4c,\,5/2-3c]$. Since 1c>34c1-c>3-4c and 5/23c<1/25/2-3c<1/2, intersection with $J$ gives
$J_{3}=[1-c,\,5/2-3c],\qquad|J_{3}|=(5/2-3c)-(1-c)=3/2-2c=\tfrac{3-2\sqrt{2}}{2}.$
For $\theta\in J\setminus J_{3}=(5/2-3c,\,1/2]$, the return time exceeds $3$. The next iterate $R_{c}^{4}\theta=\theta+4c-3$ lies in $J$ when $\theta\in[4-5c,\,7/2-4c]$. Intersection with (5/23c, 1/2](5/2-3c,\,1/2] gives
$J_{4}=(4-5c,\,1/2],\qquad|J_{4}|=1/2-(4-5c)=5c-7/2=\tfrac{5\sqrt{2}-7}{2}.$
The remainder $J_{7}=(5/2-3c,\,4-5c]=J\setminus(J_{3}\sqcup J_{4})$ has length
$|J_{7}|=(4-5c)-(5/2-3c)=3/2-2c=\tfrac{3-2\sqrt{2}}{2},$
and $R_{c}^{7}\theta\in J$ for every $\theta\in J_{7}$ by direct calculation. The three-distance theorem [11, 10] rules out return times other than $3$, 44, 77, so $J=J_{3}\sqcup J_{4}\sqcup J_{7}$ up to the two measure-zero endpoints.
The orbit $(\theta_{n-1})_{n\geq 2}$ equidistributes in $[0,1)$ by Weyl’s theorem, so each $J_{i}$ is visited with density $|J_{i}|$, and the frequency of gap value $i$ among defect positions is
$f_{i}=\frac{|J_{i}|}{|J|}.$
Substituting the lengths and using $(\sqrt{2}-1)(\sqrt{2}+1)=1$ to rationalise yields
$f_{3}=f_{7}=\sqrt{2}-1,\qquad f_{4}=3-2\sqrt{2}.$
The average gap is
$3f_{3}+4f_{4}+7f_{7}=3(\sqrt{2}{-}1)+4(3{-}2\sqrt{2})+7(\sqrt{2}{-}1)=2\sqrt{2}+2=2(\sqrt{2}{+}1),$
which equals the reciprocal of the density $\rho(\mathcal{D})=(\sqrt{2}-1)/2$ established in Theorem 24(b). ∎
@@
The gap sequence of $\mathcal{D}$ belongs to the minimal subshift generated by the primitive substitution
$\sigma\colon\quad 3\mapsto 3,4,\qquad 4\mapsto 3,7,\qquad 7\mapsto 3,7,7.$ (34)
The substitution has incidence matrix
$M=\begin{pmatrix}1&1&1\\ 1&0&0\\ 0&1&2\end{pmatrix}$,
characteristic polynomial $\lambda^{3}-3\lambda^{2}+\lambda+1=(\lambda-1)(\lambda^{2}-2\lambda-1)$, and Perron eigenvalue $1+\sqrt{2}$ (the silver ratio). Since $M^{2}$ has all positive entries, $\sigma$ is primitive of index $2$, and the Perron eigenvector gives letter frequencies matching Proposition 25.
The conjecture is supported by numerical evidence (verified for $k\leq 10^{5}$) and by the following natural candidate proof, which we leave for future work to complete.
Sketch of a Rauzy induction argument. Using the sub-intervals $J_{3}$, $J_{4}$, $J_{7}$ computed in the proof of Proposition 25, the first-return map $R_{c}^{\mathrm{ret}}\colon J\to J$ permutes the three pieces. Direct inspection of the images $R_{c}^{3}(J_{3})$, $R_{c}^{4}(J_{4})$, $R_{c}^{7}(J_{7})$ inside $J$ suggests the transitions
$J_{3}\to J_{3}J_{4},\qquad J_{4}\to J_{3}J_{7},\qquad J_{7}\to J_{3}J_{7}J_{7},$
which transpose to (34). A complete proof would require the full Rauzy induction machinery on three letters, see [7, Chapter 6].
Remark 27.
The three gap values $3,4,7$ are exactly $q_{2}$, $q_{3}-q_{2}$, and $q_{3}$ in the Pell sequence, and the conjectural Perron eigenvalue $1+\sqrt{2}$ of $\sigma$ is the dominant root of the Pell recurrence $X_{n+1}=2X_{n}+X_{n-1}$. Both facts instantiate the Pell–Ostrowski structure described in Section 4.5.